EP014C

PI: van Paradijs
Title: Sco X1

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: A
Obs time (UT): EU 03:30(13/02)-09:30(13/02)
Stations: EbJbMcNtWb
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (T. Venturi at Thu Feb 13 05:03:15 EST 1997) Success
The experiment ran smoothly without apparent problems.

Effelsberg (dgraham at Thu Feb 13 07:24:26 EST 1997) Minor failures
Ok except for heavy rain at about 0900.

Jodrell (Alastair Gunn at Fri Feb 21 03:22:56 EST 1997) Failure
POLARISATIONS ARE CROSSED. Therefore no useful data.

Calibration (Fredrik T Rantakyro at Thu May 8 11:25:36 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name ep14c.antab